quadratic equations have methods of solving
help me to solve by completing the square
4x^2+4x-15=0

Answer by bella18(6)

First you need to add 15 on both sides of the equation
4x^2+4x=15
Then divide by 4 to have 1 as leading coefficient
x^2+x=15/4
then find a value that will make the left side of the equation a perfect square. To do that, divide 1 by 2 then square it. Therefore you need 1/4 to add on both sides:
x^2+x+1/4 = 15/4 +1/4
(x+1/2)^2 =4
then get the square root of each side:
x+1/2 = +-2
x=-1/2+2= 3/2
x=-1/2-2=-5/2
therefore the solutions are 3/2 and -5/2.